OpenStudy (anonymous):

what is the inverse function of F(x)=X^5

OpenStudy (aravindg):

an inverse function to f(x) does the opposite task f(x) perfroms what will be that here @jerr1996

OpenStudy (anonymous):

i'm confused :/ would the inverse function be\[ ^5 \sqrt{x}\] ??

OpenStudy (aravindg):

yep !!

OpenStudy (aravindg):

you are right

OpenStudy (anonymous):

thank you! :)

OpenStudy (aravindg):

yw : )

OpenStudy (anonymous):

\[ F(x)=x^{2} + 5, x \ge0; G(x)=\sqrt{x-5}\] is that the inverse?

OpenStudy (klimenkov):

Yes, if you are confused try to sheck the identity: \(f(f^{-1}(x))=x\). For example: \(f(x)=x^5, f^{-1}(x)=\sqrt[5]{x},f(f^{-1}(x))=(\sqrt[5]{x})^5=x\).